OpenStudy (mhchen):

INSANE IMPOSSIBLE QUESTION 100%?!? WORLD RECORD GLITCHED POST!!! A hyperbola of x^2 - y^2 = 1 has 2 asymptotes. If we were to apply a linear transformation to the equation, would the asymptote transform the same way? T(x,y) = (3x-2,4y+1) Asymptot: y = x and y = -x

OpenStudy (fluttershyk):

Sketch those lines slopes of +/- 1 center axis halfway between 1 and 5 so at y = 3 the lines cross at (2,3), so that is center of hyperbola so (x-2)^2/a^2 - (y-3)^2/b^2 = 1 slopes of asymptotes = +/- b/a so b = a (x-2)^2/a^2 - (y-3)^2/a^2 = 1 put that point in now (4-2)^2/a^2 - (3-3)^2/a^2 = 1 a^2 = 4 a = 2 (x-2)^2/4 - (y-3)^2/4 = 1

OpenStudy (mhchen):

Thanx so much

OpenStudy (fluttershyk):

NP @mhchen